# From non-equilibrium Green functions to Lattice Wigner: A toy model for quantum nanofluidics simulations

### v1

@@ VI.2 Forcing terms
delivers $S^{q}_{0}=0$, $S^{q}_{1}=0$, $S^{q}_{2}=0$, $S^{q}_{3}=6\rho F_{3}$, $S^{q}_{4}=24JF_{3}$, where $F_{3}(x)=-\partial_{x}^{3}V$. Note that the quantum force does not act directly upon the first three moments, namely density, current and energy, although it can affect them through the gradients of the moments of order three and four. Also, since $F_{3}(x)$ stems from a third order derivative in space of the external potential, the quantum force is most effective on the short scales. For a potential lenghtscale $\delta$, the ratio of the quantum force to the classical one scales exactly like $q^{2}$.
@@ VII Numerical results
To inspect the effect of the quantum fluctuations, we study the relaxation of a one-dimensional fluid in the presence of an external potential.
The latter has the following periodic form
$V(x)=-\frac{V_{0}}{k_{n}}cos(k_{n}x),$ (34)
@@
where $k_{n}=\frac{\pi}{L}n_{w}$ and $n_{w}$ is the wavenumber. This leads to $F_{1}(x)=V_{0}sin(k_{n}x)$ and $F_{3}(x)=-V_{0}k_{n}^{2}sin(k_{n}x)$. The simulations are initialized with an inhomogeneous density distribution following a Gaussian profile and run for approximately $10^{4}$ time steps. If not stated otherwise, we study this system for two values of $n_{w}$, i.e. $n_{w}=8$ (low frequency regime) and $n_{w}=32$ (high frequency regime), two values of $\omega$, i.e. $\omega=0.5$ and $\omega=1$, $V_{0}=10^{-4}$ and $L_{x}=128$ in the absence and presence of the quantum force $F_{3}$. Also, the lattice spacings are set to $\Delta x=1$ and $\Delta t=1$ which would approximately correspond to $1$ nm and $1$ ps in real units. This leads to a channel length of roughly $128$ nm and an experiment lasting for $\sim 10$ ns. If we take $\lambda_{F}\sim 5\Delta x$ and $\delta\sim L/n_{w}$, we have $q\sim 0.3$ for $n_{w}=8$ and $q\sim 1.25$ for $n_{w}=32$, thus quantum effects are expected to become visible at high wavenumbers.
VII.1 Low wavenumber regime
@@
In Fig.3 we show the time evolution of the five power moments $P_{k}$ of classical and quantum distributions for $n_{w}=8$ and $\omega=1$ (setting the numerical viscosity to $\nu=0.5$), where the initial Gaussian profile of the density is centered at $L/2$ with a standard deviation $\sigma=4$. Classical profiles are obtained by setting $F_{3}=0$, while quantum ones include $F_{3}$. At $t=0$, all moments except $P_{0}$ are zero. Our results show that the first three moments, $P_{0}$, $P_{1}$ and $P_{2}$, are not affected by quantum forces to any appreciable extent, not even through gradients of higher order moments. Both classical and quantum distributions of $P_{0}$ and $P_{2}$ gradually relax towards an almost flat profile with values slightly larger than $1$ (Fig.3a), while $P_{1}$ displays a wave-like symmetric profile (Fig.3,b,c). The first moment $P_{1}$ is positive for $x<L/2$ and negative elsewhere, with fixed zeroes at the boundaries and at $L/2$ (i.e. where density gradients are constant), while maximum and minimum (corresponding to the inflection points of the density profile) gradually shift towards lower values, until the current vanishes everywhere.
@@
Quantum effects are found to very mildly affect only the moment $P_{3}$ (see Fig.3d,f), whose quantum distribution slightly deviates from the classical one, which displays a wave-like symmetric profile overall akin to $P_{1}$. The distribution $P_{4}$ follows the typical behavior of the even moments and is basically unaffected by any quantum deviations (Fig.3e). The different response of the moments $P_{3}$ and $P_{4}$ to the quantum force depends on the fact that, for the even moments, the effect of such a force is masked by the thermal fluctuations which, on the contrary, vanish at equilibrium for the odd moments. Note also that amplitude and frequency of both distributions at late times (Fig.3g) remain essentially consistent with the values of amplitude $V_{0}$ and wavenumber $n_{w}$ set by the periodic potential $V(x)$.
@@
Deviations from the classical distribution can be approximately quantified in terms of the percentage error $\Delta_{k}=100\times\frac{|P_{k}^{cl}-P_{k}^{q}|}{P_{k}^{cl}}$, where $P_{k}^{cl}$ and $P_{k}^{q}$ stand for classical and quantum distributions. As previously mentioned, $\Delta_{k}$ is negligible for all moments except $P_{3}$, where the highest value is found around $7\%$.
@@
Figure 3: Power moments $P_{0}$, $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ at simulation time $t=0$, $250$, $500$, $750$, $1000$, $5000$ for $n_{w}=8$ and $\omega=1$. The classical (”cl”) and quantum (”q”) distributions of the first three moments (a,b,c) are basically indistinguishable. While $P_{0}$ and $P_{2}$ relax in a similar manner starting from a Gaussian profile, $P_{1}$ exhibits an intermediate wave-like behavior, which progressively flattens towards a uniform profile. The effect of the quantum force is slightly visible for $P_{3}$ where the quantum distribution barely deviates from the classical one, as highlighted in the late-time profile shown in (f). Such effects are essentially indiscernible for $P_{4}$ (e and g).
VII.2 High wavenumber regime
@@
In Fig.4 and Fig.5 we show, for example, the time evolution of the even and odd power moments for $n_{w}=32$ and $\omega=1$. While the time behavior (Fig.4a,b,c and Fig.5a,b) is overall akin to the one observed for lower values of $n_{w}$, the effect of the quantum force is clearly visible on the profiles of all power moments (as shown in Fig.4d,e,f and Fig.5c,d). Indeed, although the quantum force enters explicitly only the moments $M_{3}$ and $M_{4}$ in Eqs.(27-28) (and thus $P_{3}$ and $P_{4}$), its effect actually conditions lower moment through spatial derivatives (for instance $\partial_{x}M_{3}$ in Eq.(26)) resulting from the hierarchical structure of the equations of motion.
@@
In the present problem, the quantum force manifests through a change of amplitude of the distributions, which display a wave-like behavior with a well-defined frequency (set by $n_{w}$). Considering, for example, the late-time profile of $P_{4}$ in Fig.4f, the quantum force amplifies the classical signal, whose minima correspond to the maxima of the quantum distribution (leading to an apparent phase-shift effect). This happens because $F_{1}$ and $F_{3}$ carry opposite signs and different amplitudes, where $F_{1}\propto V_{0}$ and $F_{3}\propto-V_{0}k^{2}_{n}$. Lower moments generally show similar features, although maxima and minima of classical and quantum distribution are found to align (such as in $P_{0}$, $P_{1}$ and $P_{3}$) and the amplitude of the quantum signal can decrease (as in $P_{2}$). Finally, the odd moments show values of $\Delta_{k}$ considerably higher than the even ones. Specifically, we find that the highest values are $\Delta_{1}\simeq 30\%$ and $\Delta_{3}\simeq 50\%$ (a difference arising essentially because the effect of the quantum force is mitigated for lower moments), while $\Delta_{0,2,4}$ are lower than $1\%$, although not negligible.
@@
It is also of interest to understand whether modifying the viscosity alters such a picture. In Fig.5 we show the time evolution of the power moments for $\omega=0.5$, which sets a numerical viscosity $\eta=1.5$. Note that, although $20<\omega<$ 2, the AdS-CFT minimum viscosity bound would impose a lower upper bound, which would be safely set at $\omega\simeq 1.5$. Once again quantum effects condition all moments, in a way overall akin to the scenario observed for the larger values of $\omega$ and with similar values of $\Delta_{k}$. Note that increasing the viscosity leads to a decrease (both in the classical and quantum distributions) of the amplitudes of the odd moments $P_{1}$ ad $P_{3}$ (corresponding to current and energy flux respectively) with respect to $\omega=1$ and to fully equilibrated (without any modulation-like effect) late-time profiles, essentially because larger values of $\eta$ entail a faster equilibration.
@@
Figure 4: Even power moments $P_{0}$ (a), $P_{2}$ (b), $P_{4}$ (c) at $t=0$, $250$, $500$, $750$, $1000$, $5000$ for $n_{w}=32$ and $\omega=1$. Figures (d), (e), (f), highlight the moment profiles at $t=5000$. The time evolution is overall akin to that observed for $n_{w}=8$. However, here the quantum force generally induces an amplitude change, either larger or smaller than that of the classical profile, as shown in the late-time profiles (d,e,f). The misalignment of maxima and minima between the distributions (for example in $P_{3}$ and $P_{4}$) results from the signed amplitude of the forces, $F_{1}\propto V_{0}$ and $F_{3}\propto-V_{0}k^{2}_{n}$.
@@
Figure 5: Odd power moments $P_{1}$ (a), $P_{3}$ (b), at $t=0$, $250$, $500$, $750$, $1000$, $5000$ for $n_{w}=32$ and $\omega=1$. Figures (d), (e), (f), highlight the moment profiles at $t=5000$. The time evolution of the distributions is similar to the one observed for lower wavenumbers, although here larger values of $n_{w}$ induce slightly wider undulations. The quantum force yields, once again, a change of amplitude with respect to the classical signal (c,d).
@@
Figure 6: Power moments $P_{0}$ (a), $P_{1}$ (b), $P_{2}$ (c), $P_{3}$ (d) and $P_{4}$ (e) at $t=0$, $250$, $500$, $750$, $1000$, $5000$ for $n_{w}=32$ and $\omega=0.5$. Figures (f), (g), (h), (i) and (j) highlight the moment profiles at $t=5000$. Unlike the previous case, here lower values of $\omega$ mainly affect odd moments, which display an undulated profile with smaller amplitude. At late times, both $P_{1}$ and $P_{3}$ do not show appreciable modulations; this is because, at larger viscosity, the fluid equilibrates on a shorter time scale.
@@ VIII Conclusions
Summarizing, we have presented a mathematical derivation of a high-order Boltzmann-Wigner lattice kinetic equation starting from non-equilibrium Green function formulation of quantum non-equilibrium transport phenomena. Simulations of a minimal D1Q5 lattice with third order quantum forcing terms for the case of a periodic potential indicate that in the semiclassical regime ($q<1$) the lowest hydrodynamic modes are well protected against quantum interference effects as long as the wavenumber $n_{w}$ (controlling the characteristic lenghtscale $\delta$) is sufficiently low. In actual practice, it is reasonable to assume that the length scale of the effective potential be significantly larger than the relevant Fermi wavelength, namely $k_{F}\delta>1$, where $k_{F}$ is the modulus of the Fermi wavevector. Of course, such an assumption needs to be checked on a case-by-case basis, but the fact remains that the lowest order moments (i.e. density and current) can only be affected by quantum interference effects on condition of strong coupling with classical non-equilibrium effects carried by the spatial gradient of the ”handshaking” moment $P_{2}$. This is indeed observed at larger values of $n_{w}$ (when $q>1$), where the presence of the quantum force generally yields a substantial change of amplitude of the classical signal. This is particularly relevant for odd moments where thermal fluctuations vanish at equilibrium, thus allowing quantum effects to emerge. It is therefore plausible to expect that the lattice Boltzmann-Wigner equation discussed in this paper may provide an efficient description of a variety of quantum-nanofluidic phenomena.
Data availability statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Conflict of interest
The authors have no conflicts to disclose.
IX Acknowledgements
The authors are grateful to L. Bocquet, N. Kavokine and T. Kaxiras for many valuable hints and discussions. M. L. and A. T. acknowledge the support of the Italian National Group for Mathematical Physics of INdAM (GNFM-INdAM).

### v2

@@ VI.2 Forcing terms
delivers $S^{q}_{0}=0$, $S^{q}_{1}=0$, $S^{q}_{2}=0$, $S^{q}_{3}=6\rho F_{3}$, $S^{q}_{4}=24JF_{3}$, where $F_{3}(x)=-\partial_{x}^{3}U$. Note that the quantum force does not act directly upon the first three moments, namely density, current and energy, although it can affect them through the gradients of the moments of order three and four. Also, since $F_{3}(x)$ stems from a third order derivative in space of the external potential, the quantum force is most effective on the short scales. For a potential lenghtscale $\delta$, the ratio of the quantum force to the classical one scales exactly like $q^{2}$.
@@ VII Numerical results
VII.1 Constant force
As a benchmark test, we consider a one-dimensional fluid subject to a constant force due, for example, to an external electric field E, and in the presence of a friction force $-\gamma{\bf u}$, where $\gamma$ is the friction coefficient. This is a basic test to verify the Ohm’s law, where an electric current $I$ (induced by the electric field) attains, at the steady state, a constant value depending on $\gamma$ (which would play the role of an electric resistance). Note that, since we deal with a 1d fluid, the inclusion of such a friction force is crucial for a correct modeling of dissipative effects, which are not fully captured by the sole shear viscosity.
In a one dimensional channel, the current $I$ coincides with its density $J$ (i.e. the first kinetic moment $M_{1}$) which, by definition, is equal to $\rho u$. At the steady state, the balance between friction and electric force leads to $u=\frac{qE}{m\gamma}$ (being $q$ the charge of the particle and $m$ the mass), finally yielding $I\propto\gamma^{-1}$. This is basically the behavior shown in Fig.3, where we plot the steady state values of $I$ versus the magnitude of the electric field $E$. As expected, the values of $I$ diminish for increasing $\gamma$, whose values can be computed as the inverse of the slope of each curve.
Figure 3: This plot shows the behavior of the steady state current $I$ induced by a constant electric field $E$. For increasing values of $\gamma$ the slope decreases, accurately following the equation $I=\frac{\rho qE}{\gamma}$. In our simulations $q$ is set to one. The values of $\gamma$ can be computed, from a linear fit, as the inverse of the slope.
Note that quantum fluctuations would be absent in this system, since they are expected to emerge when higher order derivatives of the potential survive. In the next section we consider precisely this scenario.
VII.2 Periodic potential
@@
To inspect the effect of the quantum fluctuations, we study the relaxation of a one-dimensional fluid which, in addition to the forces previously considered, is also subject to a periodic one, stemming from an external periodic potential.
The latter one has the following form
$U(x)=-U_{0}cos(k_{n}x),$ (34)
@@
where $k_{n}=\frac{\pi}{L}n_{w}$ and $n_{w}$ is the wavenumber. This leads to $F_{1}(x)=U_{0}k_{n}sin(k_{n}x)$ and $F_{3}(x)=-U_{0}k_{n}^{3}sin(k_{n}x)$. This choice ensures the existence of high order derivatives necessary to model quantum fluctuations, while the presence of a constant force plus a frictional one guarantee a nonzero steady state current and the inclusion of dissipative effects, respectively. The simulations are initialized with an inhomogeneous density distribution following a Gaussian profile and run for approximately $10^{4}$ time steps. If not stated otherwise, we study this system for two values of $n_{w}$, i.e. $n_{w}=8$ (low frequency regime) and $n_{w}=32$ (high frequency regime), two values of $\omega$, i.e. $\omega=0.5$ and $\omega=1$, $E=10^{-6}$, $\gamma=10^{-3}$, $U_{0}=10^{-3}$ and $L_{x}=128$ in the absence and presence of the quantum force $F_{3}$. Also, the lattice spacings are set to $\Delta x=1$ and $\Delta t=1$ which would approximately correspond to $1$ nm and $1$ ps in real units. This leads to a channel length of roughly $128$ nm and an experiment lasting for $\sim 10$ ns. If we take $\lambda_{F}\sim 5\Delta x$ and $\delta\sim L/n_{w}$, we have $q\sim 0.3$ for $n_{w}=8$ and $q\sim 1.25$ for $n_{w}=32$, thus quantum effects are expected to become visible at high wavenumbers.
VII.3 Low wavenumber regime
@@
In Fig.4 we show the time evolution of the five power moments $P_{k}$ of classical and quantum distributions for $n_{w}=8$ and $\omega=1$ (setting the numerical viscosity to $\nu=0.5$), where the initial Gaussian profile of the density is centered at $L/2$ with a standard deviation $\sigma=4$. Classical profiles are obtained by setting $F_{3}=0$, while quantum ones include $F_{3}$. At $t=0$, all moments except $P_{0}$ are zero. Our results show that the first three moments, $P_{0}$, $P_{1}$ and $P_{2}$, are not affected by quantum forces to any appreciable extent, not even through gradients of higher order moments. Both classical and quantum distributions of $P_{0}$ and $P_{2}$ gradually relax towards an almost flat profile with values slightly larger than $1$ (Fig.4a,c), while $P_{1}$ displays a wave-like symmetric profile (Fig.4,b). The first moment $P_{1}$ is positive for $x<L/2$ and negative elsewhere, with fixed zeroes at the boundaries and at $L/2$ (i.e. where density gradients are constant), while maximum and minimum (corresponding to the inflection points of the density profile) gradually shift towards lower values, until the current vanishes everywhere.
@@
Quantum effects are found to very mildly affect only the moment $P_{3}$ (see Fig.4d,f), whose quantum distribution slightly deviates from the classical one, which displays a wave-like symmetric profile overall akin to $P_{1}$. The distribution $P_{4}$ follows the typical behavior of the even moments and is basically unaffected by any quantum deviations (Fig.4e,g). The different response of the moments $P_{3}$ and $P_{4}$ to the quantum force depends on the fact that the effect of such a force for the even moments is masked by the thermal fluctuations which, on the contrary, vanish at equilibrium (i.e. when $u\simeq 0$) for the odd moments (see the explicit expressions of the power moments in Sec.VI.1). Note also that amplitude and frequency of both distributions at late times (Fig.4e,g) remain essentially consistent with the values of amplitude $U_{0}$ and wavenumber $n_{w}$ set by the periodic potential $U(x)$.
@@
Deviations from the classical distribution can be approximately quantified in terms of the percentage error $\Delta_{k}=100\times\frac{|P_{k}^{cl}-P_{k}^{q}|}{P_{k}^{cl}}$, where $P_{k}^{cl}$ and $P_{k}^{q}$ stand for classical and quantum distributions. As previously mentioned, $\Delta_{k}$ is negligible for all moments except $P_{3}$, where the highest value is found around $1\%$.
@@
Figure 4: Power moments $P_{0}$, $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ at simulation time $t=0$, $250$, $500$, $750$, $1000$, $10000$ for $n_{w}=8$ and $\omega=1$. The classical (”cl”) and quantum (”q”) distributions of the first three moments (a,b,c) are basically indistinguishable. While $P_{0}$, $P_{2}$ and $P_{4}$ relax in a similar manner starting from a Gaussian profile, $P_{1}$ exhibits an intermediate wave-like behavior, which progressively turns into a sinusoidal profile fluctuating around $10^{-3}$, i.e. the steady state value of the current. The effect of the quantum force is slightly visible for $P_{3}$ where the quantum distribution barely deviates from the classical one, as highlighted in the late-time profile shown in (f). Such effects are essentially indiscernible for $P_{4}$ (e and g).
VII.4 High wavenumber regime
@@
In Fig.5 and Fig.6 we show, for example, the time evolution of the even and odd power moments for $n_{w}=32$ and $\omega=1$. While the time behavior (Fig.5a,b,c and Fig.6a,b) is overall akin to the one observed for lower values of $n_{w}$ (except that here a higher wavenumber induces larger undulations), the effect of the quantum force is clearly visible on the profiles of all power moments (as shown in the late time configurations of Fig.5d,e,f and Fig.6c,d). Indeed, although the quantum force enters explicitly only the moments $M_{3}$ and $M_{4}$ in Eqs.(27-28) (and thus $P_{3}$ and $P_{4}$), its effect actually conditions lower moment through spatial derivatives (for instance $\partial_{x}M_{3}$ in Eq.(26)) resulting from the hierarchical structure of the equations of motion.
@@
In the present problem, the quantum force manifests through a change of amplitude of the distributions, which display a wave-like behavior with a well-defined frequency (set by $n_{w}$). Considering, for example, the late-time profiles of $P_{2}$ and $P_{4}$ in Fig.5e-f, the quantum force slightly amplifies the classical signal, while the opposite holds for $P_{0}$, where the minima are found to correspond to the maxima of the quantum distribution, leading to an apparent phase-shift effect. This actually happens because $F_{1}$ and $F_{3}$ carry opposite signs and different amplitudes, where $F_{1}\propto k_{n}U_{0}$ and $F_{3}\propto-U_{0}k^{3}_{n}$. Odd moments generally show similar features, although the amplitude of the quantum signal can considerably change, either increase, as in $P_{1}$, or substantially decrease as in $P_{3}$ (see Fig.6c,d). This is caused, once again, by the absence of thermal fluctuations at equilibrium for the odd moments, thus exposing such moments to quantum effects. In this respect, the odd moments show values of $\Delta_{k}$ considerably higher than the even ones. Specifically, we find that the highest values are $\Delta_{1}\simeq 13\%$ and $\Delta_{3}\simeq 30\%$ (a difference arising essentially because the quantum force is mitigated at lower moments), while $\Delta_{0,2,4}$ are lower than $1\%$, although not negligible.
@@
It is also of interest to understand whether modifying the shear viscosity alters such a picture. In Fig.7 we show the time evolution of the power moments for $\omega=0.5$, which sets a numerical viscosity $\eta=1.5$. Note that, although $20<\omega<$ 2, the AdS-CFT minimum viscosity bound would impose a lower upper bound, which would be safely set at $\omega\simeq 1.5$. Once again quantum effects condition all moments, in a way overall akin to the scenario observed for the larger values of $\omega$ and with similar values of $\Delta_{k}$. Note that increasing the viscosity leads to a slight decrease (both in the classical and quantum distributions) of the amplitudes of the odd moments $P_{1}$ ad $P_{3}$ (corresponding to current and energy flux respectively) with respect to $\omega=1$, essentially because larger values of $\eta$ entail a higher dissipation.
@@
Figure 5: Even power moments $P_{0}$ (a), $P_{2}$ (b), $P_{4}$ (c) at $t=0$, $250$, $500$, $750$, $1000$, $10000$ for $n_{w}=32$ and $\omega=1$. Figures (d), (e), (f), highlight the moment profiles at $t=10000$. The time evolution is overall akin to that observed for $n_{w}=8$. However, here the quantum force generally induces an amplitude change, either larger or smaller than that of the classical profile, as shown in the late-time profiles (d,e,f). The misalignment of maxima and minima between the distributions (for example in $P_{3}$ and $P_{4}$) results from the signed amplitude of the forces, $F_{1}\propto U_{0}k_{n}$ and $F_{3}\propto-U_{0}k^{3}_{n}$.
@@
Figure 6: Odd power moments $P_{1}$ (a), $P_{3}$ (b), at $t=0$, $250$, $500$, $750$, $1000$, $10000$ for $n_{w}=32$ and $\omega=1$. Figures (d), (e), (f), highlight the moment profiles at $t=10000$. The time evolution of the distributions is similar to the one observed for lower wavenumbers, although here larger values of $n_{w}$ induce wider undulations. The quantum force yields, once again, a change of amplitude with respect to the classical signal (c,d).
@@
Figure 7: Power moments $P_{0}$ (a), $P_{1}$ (b), $P_{2}$ (c), $P_{3}$ (d) and $P_{4}$ (e) at $t=0$, $250$, $500$, $750$, $1000$, $10000$ for $n_{w}=32$ and $\omega=0.5$. Figures (f), (g), (h), (i) and (j) highlight the moment profiles at $t=10000$. Unlike the previous case, here lower values of $\omega$ mainly affect odd moments, which display an undulated profile with smaller amplitude. At late times, both $P_{1}$ and $P_{3}$ do not show appreciable modulations; this is because, at larger viscosity, the fluid equilibrates on a shorter time scale.
@@ VIII Conclusions
Summarizing, we have presented a mathematical derivation of a high-order Boltzmann-Wigner lattice kinetic equation starting from non-equilibrium Green’s function formulation of quantum non-equilibrium transport phenomena. Simulations of a minimal D1Q5 lattice with third order quantum forcing terms for the case of a periodic potential indicate that in the semiclassical regime ($q<1$), the lowest hydrodynamic modes are well protected against quantum interference effects as long as the wavenumber $n_{w}$ (controlling the characteristic lenghtscale $\delta$) is sufficiently low. In actual practice, it is reasonable to assume that the length scale of the effective potential be significantly larger than the relevant Fermi wavelength, namely $k_{F}\delta>1$, where $k_{F}$ is the modulus of the Fermi wavevector. Of course, such an assumption needs to be checked on a case-by-case basis, but the fact remains that the lowest order moments (i.e. density and current) can only be affected by quantum interference effects on condition of strong coupling with classical non-equilibrium effects carried by the spatial gradient of the ”handshaking” moment $P_{2}$. This is indeed observed at larger values of $n_{w}$ (when $q>1$), where the presence of the quantum force generally yields a substantial change of amplitude of the classical signal. This is particularly relevant for odd moments where thermal fluctuations vanish at equilibrium, thus allowing quantum effects to emerge. It is therefore plausible to expect that the lattice Boltzmann-Wigner equation discussed in this paper may provide an efficient description of a variety of quantum-nanofluidic phenomena, although further studies are needed to investigate the interplay between external fields, dissipation and fluctuation-induced potentials. In this respect, the present work may have a limited relevance to experiments because of the use of toy potentials and the restriction to one-dimensional fluids. We plan to partially overcome these drawbacks in future works, which will be concerned with the study of two dimensional fluids (simulated on high order lattices) in the presence of more realistic potentials, such as the screened Coulomb one, and density fluctuations of the liquid.
Acknowledgements.
We thank L. Bocquet of École Normale Supérieure, N. Kavokine of EPFL and E. Kaxiras of Harvard Physics Department for many valuable hints and discussions. M. L. and A. T. thank the Italian National Group for Mathematical Physics of INdAM (GNFM-INdAM) for the support.